# HJRNO: Hamilton-Jacobi Reachability with Neural Operators

### v1

@@ 4 Results
Figure 2: Predicted BRTs along the diagonal of the hyperparameter space
To test generalization, we predict along the diagonal of the hyperparameter space. Fig. 2 illustrates the results, where predictions are made for random hyperparameter values along this diagonal corresponding to hyperparameter values not present in the training set.
Although we visualize only slices of the BRT over the $x_{1}$ (position $x$) and $x_{2}$ (position $y$) dimensions, all predictions in this section are performed over the full four-dimensional state space, including velocity $v$ and heading angle $\theta$. The results shown are obtained by randomly slicing along the $v$ and $\theta$ dimensions.
Figure 4: Random obstacle shapes
@@ 5 Discussion
The results demonstrate that HJRNO performs impressively across diverse tasks, suggesting strong potential as a foundation model for solving BRT problems.
@@
One limitation is the treatment of hyperparameters. When introducing additional hyperparameters, each value must be repeated across the input domain, increasing the size of the training dataset proportionally. This approach could become a computational bottleneck when scaling to high-dimensional hyperparameter spaces. A future research direction could be to design more efficient architectures that better handle hyperparameters, particularly those that are spatially constant, without requiring dataset duplication.

### v2

@@ 5.2 Dynamic Dubins car
Figure 5: FNO predictions along the diagonal of the hyperparameter space
Figure 6: Two Obstacles
Figure 7: Velocity-Dependent
Ground TruthPredictionError MapFNORefer to captionRefer to captionRefer to captionTNORefer to captionRefer to captionRefer to caption
Figure 8: Air3D
Trained Resolution3x Larger ResolutionFNORefer to captionRefer to captionTNORefer to captionRefer to captionCNNRefer to captionRefer to caption
Figure 9: Zero-shot super-resolution: Unlike conventional neural networks such as CNNs, our neural operators (FNO and TNO) learn mappings between function spaces, enabling zero-shot super-resolution. This allows models trained on low-resolution data to generalize directly to higher resolutions without retraining.
FNOTNOComparisonRefer to captionRefer to captionIntial ValuesRefer to captionRefer to captionGround TruthRefer to captionRefer to captionPredictionRefer to captionRefer to caption
Figure 10: Indoor Environment
Ground TruthPredictionError MapFNORefer to captionRefer to captionRefer to captionTNORefer to captionRefer to captionRefer to caption
Figure 11: Velocity-Dependent
@@ 6 Discussion
One limitation is the treatment of hyperparameters. When introducing additional hyperparameters, each value must be repeated across the hyperparameter domain, increasing the size of the training dataset proportionally. This approach could become a computational bottleneck when scaling to high-dimensional hyperparameter spaces. A future research direction could be to design more efficient approaches that better handle hyperparameters.